How many gallons of 50% antifreeze solution must be mixed with 100 gallons of 30% antifreeze to get a mixture that is 40% antifreeze?
Answer by stanbon(75887)

How many gallons of 50% antifreeze solution must be mixed with 100 gallons of 30% antifreeze to get a mixture that is 40% antifreeze?
-----------------------
Equation:
alcohol + alcohol = alcohol
0.50x + 0.30*100 = 0.40(x+100)
---
Multiply thru by 100 to get:
50x + 30*100 = 40x + 40*100
10x = 10*100
x = 100 gallons (amt. of 50% solution needed)
================================================
Cheers,
Stan H
=============